Assume you can earn 9% on the investments described below. How much money would each investment provide for you after six years? a. Invest $6,300 as a lump sum today. b. Invest $2,441 at the end of each of the next 6 years c. Invest a lump sum of $4,299 today and $1,239 at the end of each of the next 6 years d. Invest $1,111 at the end of years 1, 3, and 5
Answer #1

A. invest 6300 as lumpsom today :

there fore FV of the investment will be : 6300 (1.09)^6 = $10,566 at the end of 6 years. USING FORMULA FV= (1+R)^N

B. yearly investments are at the end of each year :

so lets consider 5 annual payments at the beginning of each year :

so PMT function is = $ 2441

n=5 years

i/y = 9%

so, the future value of the investments will be : $ 14,603

C. lump sum of $ 4299 today

so fv of this investment will be : 4299 (1.09)^ 6 = $ 7209 using formula FV=PV (1+R)^N

again 1239 at the end of each year for 6 years ,

so, PMT function = $1239

N = 5 years

i/y = 9%

now inputting these numbers in the BA 2 plus calculator we get ,

FV = $7415

TOTAL FV IS : $ 14,624( $7415 + 7209)

D. INVEST 1111 AT THE END OF 1 YEAR :

ITS VALUE HAS TO BE ACCUMULATED FOR 5 YEARS WE GET ,

1111(1.09)^5 = $1709

AT THE END OF 3 YEARS ,

1111(1.09)^3 = $ 1439

SINCE IT IS INVESTED AST THE END OF 3 YEARS,MEANING BEGINNING OF 4TH YEAR SO TOTAL NUMBER OF YEARS IT WILL BE ACCUMULATED WILL BE 4,5 AND 6 I.E 3 YEARS

AT THE END OF 5 YEARS ,

1111(1.09) = $ 1211

SO TOTAL VALUE IS :$1709 + $1439 + $1211=$ 4359